Can someone help me deduce the oxydation numbers?

Can someone help me deduce the oxydation numbers?

This is the answer :

In this reaction, the oxidation number of the carbon atom bonded to the oxygen changes from 1 in propanal to 3 in propanoic acid. Propanal is therefore oxidized to propanoic acid. The KMnO4 is reduced in this reaction.

But I got +2 for the oxydation number of the carbon in the ractants and +4 for the carbon in the products... please help!
Answer #1

We know that conversion of aldehyde to carboxylic acid results in oxidation of the carbonyl carbon

so we need to find out the oxidation number of the carbonyl carbon

-1 он H +1 1 н.с. +21.1 -2

As we can see the total charge can be calculated

Aldehyde : +2 charge due to -2 charge on oxygen and -1 charge due to +1 charge of hydrogen

So total charge = +2 -1 = +1

Acid: +2 charge due to -2 charge on oxygen and +1 charge due to -1 charge of OH.

so total charge = 2+1 = 3